### v1

# EvoSeed: Unveiling the Threat on Deep Neural Networks with Real-World Illusions

Abstract.
Deep neural networks are exploited using natural adversarial samples, which have no impact on human perception but are misclassified. Current approaches often rely on the white-box nature of deep neural networks to generate these adversarial samples or alter the distribution of adversarial samples compared to training distribution. To alleviate the limitations of current approaches, we propose EvoSeed, a novel evolutionary strategy-based search algorithmic framework to generate natural adversarial samples. Our EvoSeed framework uses auxiliary Diffusion and Classifier models to operate in a model-agnostic black-box setting. We employ CMA-ES to optimize the search for an adversarial seed vector, which, when processed by the Conditional Diffusion Model, results in an unrestricted natural adversarial sample misclassified by the Classifier Model. Experiments show that generated adversarial images are of high image quality and are transferable to different classifiers. Our approach demonstrates promise in enhancing the quality of adversarial samples using evolutionary algorithms. We hope our research opens new avenues to enhance the robustness of deep neural networks in real-world scenarios.
Deep Learning, Computer Vision, CMA-ES, Diffusion Models, Natural Adversarial Samples
Project Website can be accessed at https://shashankkotyan.github.io/EvoSeed
1. Introduction
Figure 1. Illustration of EvoSeed’s performance in finding adversarial samples under various conditions compared to Random Search (RandSeed). Our proposed EvoSeed algorithm finds high-quality Natural Adversarial Samples with a high attack success rate.
Figure 2. Illustration of the EvoSeed framework to optimize initial seed vector to generate Unconstrained adversarial samples. In this approach to generating adversarial examples, a specific class from the Cifar10 dataset is chosen as the target for the attack, denoted by condition $c$. The process begins with a random initial seed vector $z$, optimized within an $L_{\infty}$ constraint using the Covariance Matrix Adaptation Evolution Strategy (CMA-ES). A conditional diffusion model $G$ utilizes the provided condition $c$ to iteratively reduce noise from the perturbed seed vector $z^{\prime}$, resulting in image distributions $x$. The CMA-ES’s state is updated based on the fitness of the perturbed seed vector $z^{\prime}$, determined by a classifier model $F$ and the condition $c$. This iterative process continues until an adversarial seed vector $z^{\prime}$ is found, which can be used to generate an adversarial sample.
Deep Neural Networks have succeeded unprecedentedly in various visual recognition tasks. However, their performance decreases when the testing distribution differs from the training distribution, as shown by Hendrycks et al. (Hendrycks et al., 2021) and Ilyas et al. (Ilyas et al., 2019). This poses a significant challenge in developing robust deep neural networks capable of handling such shifts in distribution. Adversarial samples and adversarial attacks exploit this vulnerability by manipulating images to alter distribution compared to the original distribution.
Research by Dalvi et al. (Dalvi et al., 2004) underscores that adversarial manipulations of input data often lead to incorrect predictions from classifiers, raising serious concerns about the security and integrity of classical machine learning algorithms. This concern remains relevant, especially considering that state-of-the-art deep neural networks are highly vulnerable to adversarial attacks involving deliberately crafted perturbations to the input (Madry et al., 2018; Kotyan and Vargas, 2022).
[…]
While adversarial samples (Madry et al., 2018; Su et al., 2019; Kotyan and Vargas, 2022) expose vulnerabilities in deep neural networks, their artificial nature and reliance on constrained input data limit their real-world applicability. In contrast, the challenges become more pronounced in practical situations, where it becomes infeasible to include all potential threats within the training dataset comprehensively. This heightened complexity underscores the increased susceptibility of deep neural networks to Natural Adversarial Examples proposed by Hendrycks et al. (Hendrycks et al., 2021) and Unrestricted Adversarial Examples proposed by Song et al. (Song et al., 2018). These types of adversarial samples have gained prominence in recent years as a significant avenue in adversarial attack research, as they can make substantial alterations to images without significantly impacting human perception of their meanings and faithfulness.
In this context, we present EvoSeed, the first Evolution Strategy-based algorithmic framework designed to genearate Unrestrcited Natural Adversarial Samples as shown in Figure 2. Our algorithm requires a Conditional Diffusion Model $G$ and a Classifier Model $F$ to generate adversarial samples $x$. Specifically, it leverages the Covariance Matrix Adaptation Evolution Strategy (CMA-ES) at its core to enhance the search for adversarial seed vectors $z^{\prime}$ that can generate adversarial samples $x$. The CMA-ES fine-tunes the generation of adversarial samples through an iterative optimization process based on the Classification model outputs $F(x)$, utilizing them as fitness criteria for subsequent iterations. Ultimately, our objective is to search for an adversarial seed vector $z^{\prime}$ that, when used, causes our Conditional Diffusion Model $G$ to generate an adversarial sample $x$ misclassified by the Classifier Model $F$ and also is close to the training distribution of Conditional Diffusion Model $G$ and Classifier Model $F$, as shown in Figure 1.
Our Contributions:
Framework to Generate Natural Adversarial Samples: We propose a model-agnostic black-box algorithmic framework based on an Evolutionary Strategy titled EvoSeed to generate unrestricted natural adversarial samples. Our framework can generate natural adversarial examples using any auxiliary conditional diffusion and classifier models, as shown in Figure 2.
High-Quality Natural Adversarial Samples: Our results show that adversarial samples created using EvoSeed can also misclassify a variety of robust and non-robust classifiers alike, while being similar to the training distribution.
2. Related Work
Over the past few years, generative models such as Generative Adversarial Networks (Goodfellow et al., 2020) and Diffusion Models (Sohl-Dickstein et al., 2015) have emerged as leading tools for content creation and the precise generation of high-quality synthetic data. Several studies have employed creativity to generate Adversarial Samples; some propose the utilization of surrogate models such as (Xiao et al., 2018a; Chen et al., 2023b, a; Zhao et al., 2018; Lin et al., 2023; Jandial et al., 2019; Song et al., 2018), while other advocates the perturbation of latent representations as a mechanism for generating adversarial samples (Song et al., 2018; Zhao et al., 2018).
In the initial phases of devising natural adversarial samples, Xiao et al. (Xiao et al., 2018b) employs spatial warping transformations for their generation. Concurrently, Shamsabadi et al. (Shamsabadi et al., 2020) transforms the image into the LAB color space, producing adversarial samples imbued with natural coloration.
Song et al. (Song et al., 2018) proposes first to train an Auxiliary Classifier Generative Adversarial Network (AC-GAN) and then apply the gradient-based search to find adversarial samples under its model space. Another research proposes Adversarial GAN (AdvGan) (Xiao et al., 2018a), which removes the searching process and proposes a simple feed-forward network to generate adversarial perturbations and is further improved by Jandial et al. (Jandial et al., 2019). Similarly, Chen et al. (Chen et al., 2023b) proposes the AdvDiffuser model to add adversarial perturbation to generated images to create better adversarial samples with improved FID scores.
Yet, these approaches often have one or more limitations,
a) Some approaches rely on changing the distribution of generated images compared to the training distribution of the classifier, such as (Xiao et al., 2018b; Shamsabadi et al., 2020), or
b) Some approaches rely on the white-box nature of the classifier model to generate adversarial samples such as (Song et al., 2018; Chen et al., 2023b), or
c) Some approaches rely heavily on training models to create adversarial samples such as (Xiao et al., 2018a; Song et al., 2018; Jandial et al., 2019), or
d) Some approaches rely on generating adversarial samples for specific classifiers, such as (Xiao et al., 2018a; Jandial et al., 2019).
Thus, in contrast, we propose the EvoSeed algorithmic framework, which does not suffer from the limitations mentioned above in generating adversarial samples.
3. Background and Problem Formulation
The Diffusion Model is first proposed by Sohl-Dickstein et al. (Sohl-Dickstein et al., 2015) that can be described as a Markov chain with learned Gaussian transitions. It comprises of two primary elements: a) The forward diffusion process, and b) The reverse sampling process. The diffusion process transforms an actual distribution into a familiar straightforward random-normal distribution by incrementally introducing noise. Conversely, in the reverse sampling process, a trainable model is designed to diminish the Gaussian noise introduced by the diffusion process systematically.
Let us consider a true distribution represented as $x\in\mathbb{R}$, where $x$ can be any kind of distribution such as images (Ho et al., 2020; Dhariwal and Nichol, 2021; Ho et al., 2022; Ho and Salimans, 2022), audio (Kong et al., 2021; Huang et al., 2022a, b; Kim et al., 2022), or text (Li et al., 2022). The diffusion process is then defined as a fixed Markov chain where the approximate posterior $q$ introduces Gaussian noise to the data following a predefined schedule of variances, denoted as $\beta_{1},\beta_{2}\dots\beta_{T}$:
(1) $q(x_{1:T}|x_{0}):=\prod^{T}_{t=1}~{}q(x_{t}|x_{t-1})$
where $q(x_{t}|x_{t-1})$ is defined as,
(2) $q(x_{t}|x_{t-1}):=\mathcal{N}(x_{t};~{}\sqrt{1-\beta_{t}}\cdot x_{t-1},~{}\beta_{t}I).$
Subsequently, in the reverse process, a trainable model $p_{\theta}$ restores the diffusion process, bringing back the true distribution:
(3) $p_{\theta}(x_{0:t}):=p(x_{T})\cdot\prod^{T}_{t=1}~{}p_{\theta}(x_{t-1}|x),$
where $p_{\theta}(x_{t-1}|x)$ is defined as,
(4) $p_{\theta}(x_{t-1}|x_{t}):=\mathcal{N}\left(x_{t-1};~{}\mu_{\theta}(x_{t},t),~{}\Sigma_{\theta}(x_{t},t)\right).$
where $p_{\theta}$ incorporates both the mean $\mu_{\theta}(x_{t},t)$ and the variance $\Sigma_{\theta}(x_{t},t)$, with both being trainable models that predict the value based on the current time step and the present noise.
Furthermore, the generation process can be conditioned akin to various categories of generative models (Mirza and Osindero, 2014; Sohn et al., 2015). For instance, by integrating with text embedding models as an extra condition $c$, the conditional-based diffusion model $G_{\theta}(x_{t},c)$ creates content along the description (Ramesh et al., 2022; Saharia et al., 2022; Rombach et al., 2022; Nichol et al., 2022). This work mainly uses a conditional diffusion model to construct adversarial samples.
Unrestricted Adversarial Samples: We follow the definition from Song et al. (Song et al., 2018). Given that $\mathcal{I}$ represents a collection of images under consideration that can be categorized using one of the $K$ predefined labels. Let’s consider a testing classifier $f:\mathcal{I}\rightarrow\{1,2\dots K\}$ that can give a prediction for any image in $\mathcal{I}$. Similarly, we can consider an oracle classifier $o:O\subseteq\mathcal{I}\rightarrow\{1,2\dots K\}$ different from the testing classifier, where $O$ represents the distribution of images understood by the oracle classifier. An unrestricted adversarial sample can defined as any image inside the oracle’s domain $O$ but with a different output from the oracle classifier $o$ and testing classifier $f$. Formally defined as $x\in O$ such that $o(x)\neq f(x)$. The oracle $o$ is implicitly defined as a black box that gives ground-truth predictions. As it aligns with human assessment, the set $O$ should encompass all images perceived as realistic by humans.
Optimization on Initial Seed Vector to Generate Adversarial Samples: Let’s define a conditional diffusion model $G$ that takes an initial seed vector $z$ and a condition $c$ to generate an image $x$. Based on this, we can define the image generated by the conditional diffusion model as,
(5) $x=G(z,c)\quad\text{where}\quad z\in\mathcal{N}(\mu,\alpha^{2})$
Here, $\mu$ and $\alpha$ depend on the chosen conditional model $G$.
From the definition of the image classification problem, we can define a classifier $F$ such that $F(x)\in\mathbb{R}^{K}$ is the probabilities (confidence) for all the available $K$ labels for the image $x$. We can also define the soft label of a condition or confidence of the condition $c\in\{1,2\dots K\}$ as $F()_{c}$.
Based on the following definition, generating adversarial samples using an initial seed vector can be formulated as,
(6) $z^{\prime}=z+\eta\quad\text{such that}\quad\text{argmax}~{}[F(~{}G(z+\eta,~{}c)~{})]\neq c$
Making use of the above equation, we can formally define generating an adversarial sample as an optimization problem:
[…]
However, research by Poyuan et al. (Poyuan et al., 2023) reveals that the failure points are distributed everywhere inside the space, mostly generating images that cannot be associated with the condition $c$. To navigate these failure cases, we make the problem non-trivial by searching around the space of a well-defined initial random vector $z$. We do this by imposing an $L_{\infty}$ constraint on perturbation to initial seed vector $(\eta)$, so the modified problem becomes,
[…]
where $\epsilon$ defines the search constraint around $L_{\infty}$-sphere around initial seed vector $z$.
Algorithm 1 RandSeed: Random Search on Initial Seed Vector based on Random Shift proposed by Poyuan et al. (Poyuan et al., 2023)
1:Condition $c$, Conditional Diffusion Model $G$, Classifier Model: $F$, $L_{\infty}$ constraint: $\epsilon$, number of individuals $\lambda$, number of generations $\tau$.
2:Initialize: $z\leftarrow\mathcal{N}(0,I)$
3:for gen in $\{1\dots\tau\}$ do
4:     for i in $\{1\dots\lambda\}$ do
5:         $\eta_{i}\sim\mathcal{U}(-\epsilon,\epsilon)$
6:         individual $\leftarrow z+\eta_{i}$blacktriangleright Random Shift within bounds
7:         GeneratedImage $\leftarrow G(individual,c)$
8:         logits $\leftarrow F(GeneratedImage)$
9:         if $argmax(logits)\neq c$ then
10:              return GeneratedImage
11:         end if
12:     end for
13:end for
Algorithm 2 EvoSeed - Evolution Strategy-based Search on Initial Seed Vector
1:Condition $c$, Conditional Diffusion Model $G$, Classifier Model: $F$, $L_{\infty}$ constraint: $\epsilon$, number of individuals $\lambda$, number of generations $\tau$.
2:Initialize: $z\leftarrow\mathcal{N}(0,I)$
3:Initialize: CMAES($\mu=z$, $\sigma=1$, bounds=$(-\epsilon,\epsilon)$, pop_size=$\lambda$)
4:for gen in $\{1\dots\tau\}$ do
[…]
6:     Initialise: pop_fitness $\leftarrow$ EmptyList
[…]
8:         x $\leftarrow G(z^{\prime},c)$blacktriangleright Generated Image
9:         logits $\leftarrow F(x)$
10:         if $argmax(logits)\neq c$ then
11:              return $x$
12:         end if
[…]
14:         pop_fitness.insert(fitness)
15:     end for
[…]
17:end for
4. Methodology
Here, we define two strategies for generating an adversarial sample by manipulating the initial seed vector $z$,
a) RandSeed - A naive random search based on Random Shift (Poyuan et al., 2023),
b) EvoSeed - Our proposed evolutionary strategy-based search.
4.1. RandSeed - Random Search on Initial Seed Vector to Generate Adversarial Samples
Based on the definition of generating adversarial sample as defined in Equation 6. We can define a random search based on the Random Shift of the initial seed vector proposed by Poyuan et al. (Poyuan et al., 2023). The random shift on the initial seed vector is defined as,
(9) $z^{\prime}=z+\mathcal{U}(-\epsilon,\epsilon)$
which incorporates sampling from a uniform distribution within the range of $-\epsilon$ to $\epsilon$ Using this random shift, we can search for an adversarial sample as defined in Algorithm 1.
4.2. EvoSeed - Evolution Strategy-based Search on Initial Seed Vector to Generate Adversarial Samples
Following the definition of generating adversarial sample as an optimization problem defined in Equation 8. We optimize the search for adversarial seed vector $z^{\prime}$ using Covariance Matrix Adaptation Evolution Strategy (CMA-ES) as described by Hansen and Auger (Hansen and Auger, 2011) as defined in Algorithm 2. Our algorithm contains three main components: a Conditional Diffusion Model $G$, a Classifier model $F$, and the optimizer CMA-ES. In brief, our algorithm aims to generate unbounded adversarial samples for the given classifier model $F$ by optimizing the initial seed vector $z$ to find an adversarial seed vector $z^{\prime}$. We restrict the manipulation of $z$ with an $L_{\infty}$ constraint parameterized by $\epsilon$. This constraint ensures that each value in the perturbed vector can deviate by at most $\epsilon$ in either direction from its original value. Further, to generate and evaluate the image, we also define a condition $c$, known as the Conditional Diffusion Model $G$, and the classifier model $F$.
The commencement of the algorithm involves the initialization phase, where the initial seed vector $z$ is defined as random sampling ideal normal distribution, and the optimizer CMA-ES is set up (Lines 1 and 2 of Algorithm 2).
Following the initialization, the CMA-ES optimizes the perturbation of the initial seed vector until an adversarial seed vector is found. In each generation, the perturbation $\eta$ is sampled from a multivariate normal distribution for all the individuals in the population. Subsequently, this sampled perturbation is constrained by clipping it to fit within the specified $L_{\infty}$ range, as defined by the parameter $\epsilon$ (Line 4 of Algorithm 2).
The Conditional Diffusion Model $G$ comes into play by utilizing the perturbed vector $z^{\prime}$ as its initial state by employing a denoising mechanism to refine the perturbed seed vector, thereby forming an image distribution that closely aligns with the provided conditional information $c$ (Line 7 of Algorithm 2). Here, we leverage the CIFAR-10 (Krizhevsky et al., 2009) dataset labels to condition the generative model using condition $c$, which also serves as the true label. Consequently, the generated image is processed by the Classifier Model $F$ (Line 8 of Algorithm 2). The fitness of the perturbed seed vector $z^{\prime}$ is computed using the soft label of the condition $c$ for the logits $F(x)$ calculated by the Classifier Model $F$ (Line 12 Algorithm 2). This fitness computation plays a pivotal role in evaluating the efficacy of the perturbation within the evolutionary process.
The final phase of the algorithm involves updating the state of the CMA-ES (Lines 15 Algorithm 2). This is accomplished through a series of steps encompassing the adaptation of the covariance matrix, calculating the weighted mean of the perturbed seed vectors, and adjusting the step size. These updates contribute to the iterative refinement of the perturbation to find adversarial seed vector $z^{\prime}$.
This intricate interplay between the Conditional Diffusion Model $G$, the Classifier Model $F$, and the optimizer CMA-ES is fundamental in crafting effective and unbounded adversarial samples. In essence, our methodology leverages the power of conditioning $(c)$ of the Generative Model $G$ through a dynamic interplay with Classifier Model $F$, strategically tailored to find adversarial seed vector $z^{\prime}$ to minimize the classification accuracy on the generated image, all while navigating the delicate balance between adversarial manipulation and preserving a semblance of fidelity.
Table 1. Metric values for images generated by EDM-VP and EDM-VE variants of diffusion models without any modification to initial seed vector.
Metrics EDM-VP (Karras et al., 2022) EDM-VE (Karras et al., 2022)
FID 32.19 31.15
IS 8.30 8.46
Accuracy on Standard Non-Robust (Croce et al., 2021) 95.60% 95.90%
Accuracy on Corruptions Robust (Diffenderfer et al., 2021) 95.90% 96.80%
Accuracy on $L_{2}$ Robust (Wang et al., 2023) 95.90% 95.00%
Accuracy on $L_{\infty}$ Robust (Wang et al., 2023) 93.90% 95.40%
5. Experimental Results and Analysis
Table 2. We report Attack Success Rate (ASR), Fréchet Inception Distance (FID), Inception Score (IS) and Structural Similarity (SSIM) for various diffusion and classifier models to generate adversarial samples using Random Seed (Random Search) and EvoSeed.
Models RandSeed with $\epsilon=0.1$ EvoSeed with $\epsilon=0.1$
Diffusion Classifier Image Evaluation Image Quality Image Evaluation Image Quality
ASR ()($\uparrow$)( ↑ ) FID ()($\downarrow$)( ↓ ) SSIM ()($\uparrow$)( ↑ ) IS ()($\uparrow$)( ↑ ) ASR ()($\uparrow$)( ↑ ) FID ()($\downarrow$)( ↓ ) SSIM ()($\uparrow$)( ↑ ) IS ()($\uparrow$)( ↑ )
EDM-VP (Karras et al., 2022) Standard Non-Robust (Croce et al., 2021) 57.10% 126.94 0.25 3.72 65.70% 35.16 0.90 7.87
Corruptions Robust (Diffenderfer et al., 2021) 51.50% 124.36 0.25 3.81 64.90% 37.72 0.88 7.97
$L_{2}$ Robust (Wang et al., 2023) 47.60% 125.44 0.24 3.85 81.80% 42.38 0.83 7.17
$L_{\infty}$ Robust (Wang et al., 2023) 49.60% 124.03 0.25 3.75 83.30% 42.55 0.83 6.98
EDM-VE (Karras et al., 2022) Standard Non-Robust (Croce et al., 2021) 50.20% 112.39 0.28 4.51 66.60% 34.79 0.91 7.77
Corruptions Robust (Diffenderfer et al., 2021) 42.90% 111.93 0.28 4.42 63.70% 36.74 0.88 7.95
$L_{2}$ Robust (Wang et al., 2023) 42.70% 112.51 0.28 4.40 81.00% 41.96 0.83 7.52
$L_{\infty}$ Robust (Wang et al., 2023) 40.30% 109.92 0.28 4.45 80.10% 42.51 0.82 7.14
5.1. Experimental Setup
Conditional Generation Model $G$: We choose the EDM Diffusion Model proposed by Karras et al. (Karras et al., 2022) as the generative model $G$ and compare both variants of EDM: Variation Preserving (EDM-VP) and Variation Exploading (EDM-VE). We choose the variants trained to generate images like the CIFAR-10 dataset (Krizhevsky et al., 2009).
Classifier Model $F$: We choose the models available on RobustBench Leaderboard (Croce et al., 2021) for our experiments. We specifically chose one non-robust neural network, Standard WideResNet (Zagoruyko and Komodakis, 2016), and three different robust neural networks. Specifically, we use WideResNets trained by Wang et al. (Wang et al., 2023), which are robust against $L_{2}$ adversarial attacks and $L_{\infty}$ adversarial attacks. We also use WideResNet trained by (Diffenderfer et al., 2021), which is robust against common corruptions (Hendrycks and Dietterich, 2019).
Optimizer CMA-ES: We chose to use the Vanilla Covariance Matrix Adaptation Evolution Strategy (CMA-ES) proposed by Hansen and Auger (Hansen and Auger, 2011) to optimize the initial seed vector $z$ in order to find adversarial seed vectors $z^{\prime}$ which can generate unrestricted natural adversarial samples. We initialize CMA-ES with $\mu$ with an initial seed vector and $\sigma=1$. To limit the search by CMA-ES, we also impose an $L_{\infty}$ constraint defined by the initial seed vector on the population. We further optimize for $\tau=100$ generations with a population of $\lambda=28$ individual seed vectors $z^{\prime}$. We also set up an early finish of the algorithms if we found an individual seed vector $z^{\prime}$ in the population that could misclassify the classifier model. For our experiments, we parameterize the amount of $L_{\infty}$ constraint as $epsilon$ and experiment for three different constraint values: $0.1$, $0.2$, and $0.3$.
Metrics: We perform experiments by generating $1000$ images by randomly choosing a target class for each image. We evaluate the generated images over various metrics as described below,
a) We evaluate the image by calculating the Attack Success Rate $(ASR)$ of generated images, defined as the number of images misclassified by the classifier model $F$. It defines how likely an algorithm will generate an adversarial sample.
b) We also evaluate the quality of the image by calculating two distribution-based metrics, Fréchet Inception Distance (FID) (Heusel et al., 2017), and Inception Score (IS) (Salimans et al., 2016) and by evaluating Structural Similarity Image Metric (SSIM) (Wang et al., 2004) between the images generated by non-modified initial seed vector $z$ and the optimized adversarial seed vector $z^{\prime}$.
Table 1 reports the quality of images generated using non-modified initial seed vector $z$ by the variants EDM-VP and EDM-VE $(F)$ and also reports the accuracy on different classifier models $(G)$. We observe that the images generated by the variants are high image quality and classifiable by different classifier models with over $94\%$ accuracy.
5.2. Analysis of EvoSeed over Random Search (RandSeed)
Figure 3. Exemplar adversarial samples generated using EvoSeed and RandSeed algorithms. Note that EvoSeed finds high-quality adversarial samples comparable to samples from the original CIFAR-10 dataset. In contrast, RandSeed finds low-quality, highly distorted adversarial samples with a color shift towards the pure white image.
(a) (b)                               Figure 4. Accuracy on Generated Images $x$ by the classifier model $F$ over $\tau$ generations. (a) compares the performance of EvoSeed and RandSeed, while (b) compares the performance of EvoSeed with different classifier models.
In order to compare EvoSeed with Random Search (RandSeed), Table 2 compares the performance of EvoSeed with RandSeed, a random search approach to find adversarial samples. The comparison involves evaluating EvoSeed’s potential to generate adversarial samples using various diffusion and classifier models. The results presented in Table 2 demonstrate that EvoSeed discovers more adversarial samples than Random Seed and produces higher image-quality adversarial samples. The image quality of adversarial samples is comparable to that of non-adversarial samples generated by the Conditional Diffusion Model.
The disparity in image quality between EvoSeed and RandSeed is visually depicted in Figure 3. Images generated by RandSeed exhibit low quality, marked by distortion and a noticeable color shift towards white. This suggests that employing diffusion models for a simplistic search of adversarial samples using RandSeed can yield poor-quality results. Conversely, EvoSeed generates high image-quality adversarial samples comparable to the original CIFAR-10 dataset, indicating that it can find good-quality adversarial samples without explicitly being optimized for image quality.
Efficiency in finding adversarial samples is further highlighted by EvoSeed’s superiority in utilizing $L_{2}$ Robust (Wang et al., 2023) and $L_{\infty}$ Robust (Wang et al., 2023) classifiers over Standard Non-Robust (Croce et al., 2021) and Corruptions Robust (Diffenderfer et al., 2021) classifiers. This suggests that $L_{2}$ and $L_{\infty}$ Robust models were trained on slightly shifted distributions, as evidenced by marginal changes in FID scores and IS scores of the adversarial samples. Additionally, the performance of EDM-VP and EDM-VE variants is comparable, with EDM-VP discovering slightly more adversarial samples while EDM-VE produces slightly higher image-quality adversarial samples.
To comprehend the Evolution of adversarial samples across generations, Figure 5.2 (a) illustrates how EvoSeed progressively discovers more adversarial samples with increasing generations of optimization. In contrast, RandSeed quickly finds adversarial samples within the initial generations and does not improve much later. This suggests that if the random shift can potentially generate an adversarial sample, any value of the random shift can disrupt the image generation process by the conditional diffusion model. Figure 5.2 (b) reveals the impact of the chosen classifier model $(F)$ on EvoSeed’s performance. The Corruptions Robust (Diffenderfer et al., 2021) classifier demonstrates slightly greater robustness to EvoSeed than other robust and non-robust classifier variants.
5.3. Analysis of EvoSeed over $L_{\infty}$ constraint on initial seed vector
Table 3. We report Attack Success Rate (ASR), Fréchet Inception Distance (FID), Inception Score (IS), and Structural Similarity Score (SSIM) for various diffusion and classifier models to generate adversarial samples using EvoSeed with different search constraints.
Models EvoSeed with $\epsilon=0.2$ EvoSeed with $\epsilon=0.3$
Diffusion Classifier Image Evaluation Image Quality Image Evaluation Image Quality
ASR ()($\uparrow$)( ↑ ) FID ()($\downarrow$)( ↓ ) SSIM ()($\uparrow$)( ↑ ) IS ()($\uparrow$)( ↑ ) ASR ()($\uparrow$)( ↑ ) FID ()($\downarrow$)( ↓ ) SSIM ()($\uparrow$)( ↑ ) IS ()($\uparrow$)( ↑ )
EDM-VP (Karras et al., 2022) Standard Non-Robust (Croce et al., 2021) 88.00% 38.44 0.86 7.77 92.80% 40.29 0.81 7.90
Corruptions Robust (Diffenderfer et al., 2021) 86.30% 40.04 0.83 7.76 92.40% 41.48 0.78 7.52
$L_{2}$ Robust (Wang et al., 2023) 94.60% 44.03 0.79 7.45 98.40% 44.88 0.76 7.34
$L_{\infty}$ Robust (Wang et al., 2023) 96.00% 43.67 0.79 7.69 98.50% 45.55 0.76 7.23
EDM-VE (Karras et al., 2022) Standard Non-Robust (Croce et al., 2021) 87.80% 37.05 0.86 7.59 93.70% 38.98 0.82 7.69
Corruptions Robust (Diffenderfer et al., 2021) 87.30% 39.94 0.83 7.93 92.50% 41.22 0.79 7.81
$L_{2}$ Robust (Wang et al., 2023) 95.50% 44.15 0.80 7.70 98.00% 45.30 0.76 7.54
$L_{\infty}$ Robust (Wang et al., 2023) 96.10% 44.05 0.79 7.39 98.50% 45.09 0.76 7.40
Figure 5. Accuracy on Generated Images $x$ by the classifier model $F$ over $\tau$ generations. It compares the performance of EvoSeed with different search constraints $\epsilon$.
Table 4. We report attack success rate on Standard Non-Robust Classifier (Croce et al., 2021), Corruptions Robust Classifier (Diffenderfer et al., 2021), $L_{2}$ Robust Classifier (Wang et al., 2023) and $L_{\infty}$ Robust Classifier (Wang et al., 2023) for adversarial samples generated using different diffusion and classifier models.
Models Attack Success Rate (ASR) ()normal-↑($\uparrow$)( ↑ ) on Average ASR ()normal-↑($\uparrow$)( ↑ ) on
Diffusion Classifier Standard Non-Robust (Croce et al., 2021) Corruptions Robust (Diffenderfer et al., 2021) $L_{2}$ Robust (Wang et al., 2023) $L_{\infty}$ Robust (Wang et al., 2023) 4 Classifier Models
EDM-VP (Karras et al., 2022) Standard Non-Robust (Croce et al., 2021) 92.80% 17.10% 14.10% 19.40% 35.85%
Corruptions Robust (Diffenderfer et al., 2021) 41.80% 92.40% 28.20% 34.70% 49.28%
$L_{2}$ Robust (Wang et al., 2023) 33.60% 39.40% 98.40% 68.70% 60.03%
$L_{\infty}$ Robust (Wang et al., 2023) 27.70% 27.10% 35.90% 98.50% 47.30%
EDM-VE (Karras et al., 2022) Standard Non-Robust (Croce et al., 2021) 93.70% 15.80% 14.30% 18.60% 35.60%
Corruptions Robust (Diffenderfer et al., 2021) 43.50% 92.50% 29.50% 36.10% 50.40%
$L_{2}$ Robust (Wang et al., 2023) 34.20% 36.00% 98.00% 68.10% 59.08%
$L_{\infty}$ Robust (Wang et al., 2023) 27.60% 26.10% 37.50% 98.50% 47.43%
(a) (b)                               Figure 6. Confusion Matrix of Generated Adversarial samples using EvoSeed with $\epsilon$ = 0.3 and (a) Corruptions Robust Model (Diffenderfer et al., 2021), (b) $L_{2}$ Robust Model (Wang et al., 2023).
To enhance the success rate of attacks by EvoSeed, we relax the constraint on the $L_{\infty}$ bound $\epsilon$ to expand the search space of CMA-ES. The performance of EvoSeed under various search constraints $\epsilon$ applied to the initial search vector is compared in Table 3 to identify optimal conditions for finding adversarial samples. The results in Table 3 indicate an improvement in EvoSeed’s performance, leading to the discovery of more adversarial samples, albeit with a slight compromise in image quality. Specifically, when employing an $\epsilon=0.3$, EvoSeed successfully identifies over $92\%$ of adversarial samples, regardless of the diffusion and classifier models utilized.
To comprehend the progression of adversarial samples across generations, Figure 5 illustrates how EvoSeed, under different search constraints, progressively discovers more adversarial samples as optimization generations increase.
Notably, as the search constraint is relaxed, the graph depicting EvoSeed’s performance becomes steeper, indicating a larger pool of potential adversarial seed vectors. The increased search space makes the identification of adversarial seed vectors more accessible.
6. Extended Analysis of Generated Natural Adversarial Samples
6.1. Analysis of Transferability of Generated Adversarial Samples
To assess the quality of adversarial samples, we evaluated the Attack Success Rate (ASR) across various classifier models. The transferability of adversarial samples generated under different conditions was examined, and the results are presented in Table 4, categorized by Standard Non-Robust (Croce et al., 2021), Corruptions Robust (Diffenderfer et al., 2021), $L_{2}$, and $L_{\infty}$ Robust (Wang et al., 2023) classifiers.
Analysis of Table 4 reveals that using the $L_{2}$ Robust classifier yields the highest quality adversarial samples, with approximately $60\%$ transferability across various scenarios. It is noteworthy that adversarial samples generated with the $L_{2}$ Robust classifier can also be misclassified by the $L_{\infty}$ Robust classifier, achieving an ASR of $68\%$.
We also note that adversarial samples generated by Standard Non-Robust (Croce et al., 2021) classifier have the least transferability, indirectly suggesting that the distribution of adversarial samples is closer to the original dataset as reported in Table 3. We also note that Corruptions Robust (Diffenderfer et al., 2021), $L_{2}$ Robust (Wang et al., 2023) classifiers, are the most robust models to EvoSeed algorithm with less than $43\%$ and $44\%$ ASR across different scenarios with adversarial samples generated using different diffusion and classifier models.
6.2. Class-wise Analysis of Generated Adversarial Samples
We examine the confusion matrix of adversarial samples to gain insights into the distribution of adversarial samples across conditioned labels $c$ and predicted labels by the classifier model $F$. This analysis is performed for adversarial samples generated using the Corruptions Robust (Diffenderfer et al., 2021) and $L_{2}$ Robust classifiers, as depicted in Figure 5.3.
The patterns in predicted labels are observed consistently across different classifiers through the insights provided by Figure Figure 5.3. For instance, images labeled as "Ship" predominantly tend to be predicted as "Airplane" by the classifiers. Furthermore, there is a consistent trend of confusion between specific pairs of classes, where images labeled as "Bird" are commonly misclassified as "Truck," and vice versa. Similarly, images labeled as "Cat" often experience misclassification as "Dog," and vice versa.
Significantly, these identified patterns persist irrespective of the particular classifier employed. This suggests that these inherent adversarial patterns exist within the training distribution of diffusion and classifier models, which EvoSeed effectively leverages. The black-box nature of EvoSeed complicates understanding whether these adversarial patterns are manipulated within the Diffusion Model $G$ or the Classifier Model $F$. However, a simultaneous analysis of Table 4 and Figure 5.3 indicates that EvoSeed predominantly exploits the Classifier Model rather than the Conditional Generation Model.
This is because, if the Diffusion Models were the focus of exploitation, the images generated by the Diffusion Model for the "Cat" condition would likely consist primarily of objects of the "Dog" class, leading to perfect misclassification by other classifiers since they would classify the image as "Dog" as opposed to the conditioned variable "Cat."
7. Conclusions
This study introduces EvoSeed, first-of-a-kind evolutionary strategy-based approach for generating natural adversarial samples. Our framework employs EvoSeed within a model-agnostic black-box setup, utilizing an auxiliary Conditional Diffusion Model, a Classifier Model, and CMA-ES to produce natural adversarial examples.
Experimental results demonstrate that EvoSeed excels in discovering high-quality adversarial samples closely aligned with the original data distribution, in contrast to random search methods that often yield distorted and unrecognizable adversarial samples. Additionally, we investigate the impact of varying the search space on EvoSeed’s performance and find that expanding the search space enhances the quantity of generated adversarial samples with minimal trade-off in image quality. We anticipate that this research will pave the way for new developments in generating natural adversarial samples, providing valuable insights into the limitations of classifier robustness.